The equilibrium constant, K, for the following reaction is 3.16×10-2 at 525 K.

  • PCl5(g) <---> PCl3(g) + Cl2(g)


An equilibrium mixture of the three gases in a 18.5 L container at 525 K contains 0.232 M PCl5,   8.57×10-2 M PCl3 and 8.57×10-2 M Cl2. What will be the concentrations of the three gases once equilibrium has been reestablished, if the equilibrium mixture is compressed at constant temperature to a volume of 8.42 L?

[PCl5] = M
[PCl3] = M
[Cl2] = M
Answer #1

Answer

[PCl5] = 0.5645M

[PCl3]= 0.1336M

[Cl2] = 0.1336M

Explanation

PCl5(g) <------> PCl3(g) + Cl2(g)

After changing the volume

Initial concentration

[PCl5] = 0.232M/( 8.42L/18.5L) = 0.5097M

[PCl3] = 0.0857M/(8.42L/18.5L) = 0.1883M

[Cl2] = 0.0857M/( 8.42L/18.5L) = 0.1883M

Change in concentration

[PCl5] = -x

[PCl3] = + x

[Cl2] = +x

Equilibrium concentration

[PCl5] = 0.5097 -x

[PCl3] = 0.1883 + x

[Cl2] = 0.1883 + x

so,

(0.1883 + x)2/(0.5097 -x) = 0.0316

solving for x

x = - 0.05475

Therefore, at equilibrium

[PCl5] = 0.5097 -( - 0.05475) = 0.5645M

[PCl3] = 0.1883 + ( - 0.05475) = 0.1336M

[Cl2] = 0.1883 + ( - 0.05475) = 0.1336M